if i am traveling at 35 miles per hour, how long will it take me to travel 7 miles? also, i would like to see the formula. thanks, jerry
Answer by minerva(22)

no problem,
speed=distance/time
35=7/x
cross multiply,
35x=7
x=7/35
x=1/5 of an hour
thus ,
1/5 0f 60
60/5
=12 minutes